# Loss-Oriented Ranking for Automated Visual Prompting in LVLMs

### v2

@@ Abstract
Inspired by text prompts in large language models (LLMs), visual prompts have been explored to enhance the reasoning capabilities of large vision-language models (LVLMs). Current methods design heuristic visual prompts, such as overlaying a text-query-guided attention heatmap on the original input image. However, designing effective prompts manually is challenging and time-consuming, and it often fails to explore the benefits of different visual prompts, leading to sub-optimal performance. To this end, we propose AutoV that learns to automatically select the optimal visual prompt from various candidates based on given textual queries and the input image. To train AutoV, we develop an automatic data collection and labeling pipeline that evaluates various visual prompts with a pre-trained LVLM. We input a set of visual prompts into the LVLM and rank them according to the prediction losses generated by the model. Using the ranking as a supervision signal, we train AutoV to automatically choose the optimal visual prompt from various visual prompts for LVLMs. Experiments indicate that AutoV enhances the performance of various LVLMs across multiple image understanding tasks. For instance, LLaVA-OV with AutoV achieves accuracy gain on VizWiz, and AutoV boosts Qwen2.5-VL by on MMMU, highlighting its potential as an optimal visual prompting method.
@@ 1 Introduction
Recent advancements in large language models (LLMs) [brown2020language, achiam2023gpt, touvron2023llama, peng2023instruction, bi2024deepseek] have spurred significant progress in large vision-language models (LVLMs), which connect visual encoders with language model decoders to effectively adapt textual reasoning capabilities to visual understanding tasks [liu2023improvedllava, Qwen-VL, chen2023internvl, bai2025qwen2, li2024llava]. Therefore, visual input plays a crucial role in LVLMs, enabling precise perception and fine-grained grounding across various multimodal scenarios.
To further improve the perception and reasoning of LVLMs, visual prompting [api, lin2024training, yang2023fine, shtedritski2023does] has emerged as an important technique analogous to textual prompting in LLMs [kojima2022large, zhou2022cocoop, wei2022chain]. Visual prompts enhance model performance by adding visual annotations such as masks, colorful circles, or heatmaps, thus guiding the attention of LVLMs toward crucial image regions of interest. However, existing visual prompting methods predominantly rely on heuristically designed prompts, limiting their ability to fully exploit the inherent diversity of visual prompts. As illustrated in Figure 1, we list the benchmark-level performance of existing heuristic prompt methods (e.g., FGVP [yang2023fine], Red Circle [shtedritski2023does], and API [api]). While all three methods demonstrate effectiveness on logic-intensive MMMU tasks, both blur-based FGVP and bounding prompts adversely affect vision-centric benchmarks (e.g., VizWiz [bigham2010vizwiz]) and OCR-dependent tasks (e.g., TextVQA [singh2019towards]). This is because heuristic prompt methods fail to automatically tune hyperparameters for each benchmark or even each instance.
@@
Hence, to develop a more effective visual prompt strategy suitable for diverse vision-language tasks, it is vital to adaptively retrieve the optimal visual prompt that most significantly improves the perception and reasoning of LVLMs in response to specific queries. Motivated by this, we introduce AutoV, an innovative visual prompting retrieval framework designed to dynamically select the optimal visual prompt from a set of candidates, explicitly tailored to the textual query and input image. Specifically, we first generate various visual prompt candidates and encode their representation. This diversity enables the framework to capture distinct visual representations that may be optimal for various query-image pairs. Next, we design a lightweight ranking network that efficiently integrates visual and textual information to predict the rankings preferences over prompt candidates. Finally, the highest-ranking candidate is chosen as the input for LVLMs during inference.
To train the ranking network, we establish an automated pipeline for collecting and labeling training data, evaluating the rank of each visual prompt candidate based on the prediction losses of a pre-trained LVLM. Subsequently, the ranking network is supervised using a reward loss that differentiates between selected and rejected candidates. This strategy provides reliable supervision, allowing AutoV to attain adaptive, query-specific prompt retrieval without the need for extensive manual annotation work.
AutoV achieves notable performance improvements across various prominent vision-language models and diverse image understanding benchmarks, including fine-grained comprehension, real-world visual assistance, and complex open-ended reasoning tasks. For instance, integrating AutoV with LLaVA-OneVision [li2024llava] and InternVL2 [chen2023internvl] leads to improvements of and on VizWiz [bigham2010vizwiz], respectively. Additionally, AutoV enhances Qwen2.5-VL [Qwen-VL] on MMMU [yue2024mmmu] by , highlighting its significant advantage in visual prompting reasoning retrieval. Furthermore, AutoV generalizes well, maintaining stable performance across captioning, grounding, and classification tasks. In summary, our contributions are as follows:
@@
We introduce AutoV, a novel automatic visual prompting framework for large vision-language models that adaptively retrieves the optimal visual prompt from a diverse candidate pool, explicitly tailored to each textual query.
@@
We perform comprehensive experiments to validate the efficacy and robustness of AutoV across various multimodal models and tasks, showcasing consistent and significant improvements over existing visual prompting methods. Once trained, AutoV integrates seamlessly into various LVLMs without additional fine-tuning.
@@ 2 Related Work
Large Vision-Language Models. Recent advances in large language models [radford2019language, brown2020language, achiam2023gpt, touvron2023llama, peng2023instruction, bi2024deepseek] have catalyzed a new wave of large vision-language models [liu2023improvedllava, alayrac2022flamingo, Qwen-VL, chen2023internvl, li2024mini, li2023blip]. By coupling a high-capacity visual encoder with an LLM decoder, LVLMs transfer textual reasoning to visual understanding, and the quality of input images determines the lower bound of multimodal abilities. For instance, the popular LLaVA-OneVision [li2024llava] and Qwen2.5-VL [bai2025qwen2] adopt a higher and dynamic resolution recipe for training to enhance their fine-grained reasoning, while vision-centric scaling in InternVL2 [cai2024internlm2] enlarges the ViT [dosovitskiy2020image] backbone to boost grounding capability. However, the above methods require training times of several hundred GPU days. Here, can we further boost the existing LVLMs by optimizing the visual input itself during the inference stage? The answer is Yes. Visual prompting [sun2023alphaclip, lin2024drawandunderstand, jiang2024joint, zhang2024exploring, api, yang2023fine, shtedritski2023does] is an emerging and effective technique in this direction.
@@
Visual Prompting for LVLMs. LVLMs relying solely on textual inputs exhibited two key limitations: visual hallucinations [bai2024hallucination, huang2024visual] and language-driven biases [wu2024commit, qu2024unified]. To better mitigate these challenges, visual prompting [zhang2023vpgtrans, sun2023alphaclip, lin2024drawandunderstand, wu2024visual] has emerged as an alternative way of supplying cues in the visual modality. Compared with textual prompts, it is more concise and explicit, while allowing guidance that reaches fine-grained, pixel-level details. Early work [shtedritski2023does] shows that simply drawing a red circle around a target object steers CLIP [CLIP] attention, yielding great performance on referring-expression and keypoint-localization benchmarks. FGVP [yang2023fine] extends this idea with fine-grained prompting, replacing coarse shapes with blur-reversal prompts based on segmentation masks, further improving accuracy on RefCOCO [zhang2019referring]. AlphaCLIP [sun2023alphaclip] trains an augmented version of CLIP with an auxiliary alpha channel for indicating regions at the expense of increased training overhead. API [api] proposes attention prompting on images, overlaying text-guided attention heatmaps, boosting LVLMs on VQA, and open-ended reasoning tasks. Although previous studies have explored various visual prompts, it remains unclear which prompt is optimal for a specific image-question pair, leading to a decrease in the practicability of visual prompts. To address this limitation, we introduce our visual prompting method that learns to retrieve the optimal visual prompt for a specific image-question pair.
@@
Visual Retrieval for LVLMs. A variety of strategies have been proposed for visual retrievals, such as visual token selection [chen2024image, zhang2024sparsevlm, yang2024visionzip] for LVLMs. Existing visual prompt selection pipelines generally adhere to three main paradigms. (1) The heuristic paradigm identifies benchmark-level optimal parameters through extensive evaluation without instance-specific adaptation. (2) The random paradigm offers a baseline characterized by instance-level variation. (3) The Mixture-of-Experts (MoE) paradigm [fedus2022switch, bao2022vlmo, zong2024mova] provides the most direct form of adaptation by utilizing the GateNet specifically for visual retrieval, which is susceptible to overfitting during training.
Notably, our AutoV fundamentally differs from retrieval-augmented methods [qi2024rora, xing2025re]. Unlike them, which retrieve supplementary visual data from external databases and require model finetuning, AutoV selects the most suitable visual prompt based solely on the internal features of the current image through a lightweight ranking network. Moreover, it operates within a compact prompt pool rather than a large-scale database, yielding substantial efficiency gains.
@@ 3.3 Reward Loss Supervision
Automated Data Generation. We create customized data to train the ranking network to select the optimal visual prompt for a given LVLM. Our approach is grounded in a simple intuition: a superior visual prompt should yield lower language modeling loss when input into any trained LVLM. We optimize the ranking network using relative loss instead of a one-hot selection vector, as correctness can stem from the LVLM’s language priors without relying on any input content [xiao2024can, yu2024frame]. In contrast, the inference loss indicates the confidence of models in generating answers based on the input visual-question pair: a lower loss indicates higher confidence [guo2017calibration, zhang2024unveiling].
@@ 3.4 Robust Inference Pipeline
During the inference, we input multiple candidate visual prompts along with the query into the large vision-language model and pass them through our ranking network, as shown in Figure 2. The network outputs a reward scalar for each visual prompt, and we select the highest-scoring one. Its visual tokens are concatenated with the text embeddings and decoded by the LLM to generate the final answer.
Besides, to reduce distributional bias, we introduce a pre-filtering step: the prompt whose visual features are farthest from other candidates, as measured by cosine distance, is removed in advance, further mitigating the negative effects of suboptimal prompts and enhancing robustness.
@@ 4 Experiments
4.1 Experiment Settings
Models. We apply our AutoV to four mainstream open-source large vision-language models, i.e., LLaVA-1.5 series [liu2024visual], LLaVA-OneVision [li2024llava], Qwen2.5-VL [bai2025qwen2], and InternVL2 [chen2024expanding]. In particular, LLaVA-1.5 series build on CLIP-ViT [CLIP] and Llama2 [touvron2023llama2]. LLaVA-OneVision enhances it with SigLIP [siglip] and Qwen2 [bai2023qwen]. Qwen2.5-VL introduces a dynamic-resolution encoder, and InternVL2 integrates InternViT [chen2023internvl] with InternLM2 [cai2024internlm2].
Benchmarks. AutoV is extensively evaluated on fifteen vision-language benchmarks, covering video reasoning (CVBench [tong2024cambrian]), vision-centric tests (BLINK [fu2024blink], MMVP [tong2024eyes], MMStar [chen2024we]), comprehensive multimodal evaluation (MM-Vet [yu2023mm]), knowledge/math reasoning (MMMU [yue2024mmmu], MathVista [lu2023mathvista]), robustness and VQA (VizWiz [bigham2010vizwiz], Real-World QA [RealworldQA], TextVQA [singh2019towards], LLaVA-Wild [liu2023improvedllava]), grounding (RefCOCO+ [yu2016modeling]), object-hallucination probing (POPE [li2023evaluating]), captioning (COCO-Caption2017 [lin2014microsoft]), and multimodal classification (ImageNet-1K [deng2009imagenet]).
@@ 4.2 Main Results
| Method | MMMU | POPE | VizWiz | MMVet | VQA | LLaVA | MMStar | MathVista | Avg. |
| LLaVA-v1.5 (7B) | |||||||||
| Base Model | 36.3 | 85.9 | 50.0 | 30.6 | 58.2 | 65.4 | 33.2 | 34.2 | |
| FGVP (NIPS23) | 36.9 | 72.6 | 47.5 | 29.0 | 46.3 | 55.6 | 33.0 | 34.6 | |
| Circle (ICCV23) | 37.0 | 84.5 | 49.6 | 31.3 | 55.5 | 62.2 | 33.4 | 35.0 | |
| API (ECCV24) | 37.4 | 86.6 | 51.3 | 31.8 | 58.4 | 67.2 | 34.0 | 35.7 | |
| AutoV | 38.7 | 87.4 | 52.1 | 32.9 | 60.6 | 68.6 | 35.2 | 36.9 | |
| LLaVA-v1.5 (13B) | |||||||||
| Base Model | 36.7 | 85.9 | 53.6 | 36.1 | 61.2 | 66.6 | 32.8 | 35.0 | |
| FGVP (NIPS23) | 36.9 | 76.2 | 46.8 | 29.7 | 49.2 | 63.7 | 31.9 | 35.3 | |
| Circle (ICCV23) | 36.9 | 84.6 | 53.8 | 31.2 | 58.1 | 66.9 | 33.5 | 36.2 | |
| API (ECCV24) | 37.3 | 86.8 | 54.6 | 37.6 | 62.1 | 67.6 | 33.0 | 37.0 | |
| AutoV | 38.9 | 87.8 | 56.3 | 38.0 | 62.7 | 69.5 | 34.2 | 38.3 | |
| LLaVA-OneVision (7B) | |||||||||
| Base Model | 49.4 | 89.1 | 58.2 | 48.8 | 76.1 | 80.6 | 61.8 | 63.2 | |
| FGVP (NIPS23) | 49.6 | 82.9 | 58.9 | 49.1 | 65.4 | 70.8 | 58.8 | 63.4 | |
| Circle (ICCV23) | 50.4 | 88.2 | 60.2 | 47.2 | 77.2 | 80.8 | 62.7 | 65.4 | |
| API (ECCV24) | 50.8 | 90.3 | 66.9 | 51.5 | 77.7 | 81.9 | 63.0 | 64.9 | |
| AutoV | 54.0 | 91.0 | 68.4 | 52.9 | 78.0 | 85.3 | 64.5 | 66.8 | |
| InternVL2 (8B) | |||||||||
| Base Model | 48.6 | 87.8 | 58.6 | 46.5 | 77.0 | 74.6 | 62.0 | 58.3 | |
| FGVP (NIPS23) | 49.6 | 84.2 | 58.3 | 44.3 | 70.7 | 75.7 | 60.0 | 58.9 | |
| Circle (ICCV23) | 48.3 | 86.0 | 57.5 | 45.8 | 74.9 | 75.4 | 63.4 | 58.4 | |
| API (ECCV24) | 50.5 | 88.9 | 61.2 | 48.6 | 77.0 | 76.2 | 63.9 | 59.4 | |
| AutoV | 51.7 | 89.6 | 64.8 | 50.2 | 78.0 | 79.5 | 65.8 | 70.5 | |
| Qwen2.5-VL (7B) | |||||||||
| Base Model | 50.1 | 87.5 | 69.5 | 56.6 | 83.0 | 98.0 | 62.4 | 68.2 | |
| FGVP (NIPS23) | 49.6 | 81.6 | 70.8 | 43.1 | 78.0 | 98.2 | 63.4 | 68.0 | |
| Circle (ICCV23) | 51.6 | 84.2 | 69.2 | 44.5 | 80.5 | 97.5 | 65.1 | 69.1 | |
| API (ECCV24) | 51.0 | 87.9 | 72.1 | 58.0 | 85.3 | 100.6 | 63.0 | 67.6 | |
| AutoV | 53.9 | 89.5 | 74.4 | 59.1 | 86.1 | 102.3 | 65.6 | 70.2 | |
@@
Our results are listed in Table 1, and we compare AutoV with three previous visual prompting methods designed for LVLMs, including FGVP [yang2023fine], RedCircle [shtedritski2023does], and API [api]. The training data and implementation details can be found in the Appendix A and C, respectively. Here, we employ the above three types of visual prompts111The pool includes four attention-based visual prompts extracted from different transformer layers of CLIP in API [api], together with two additional prompts of distinct types, resulting in six images in total. as the candidate pool for AutoV to retrieve from optimally.
@@
The four main observations from the experimental results are as follows: (1) Our AutoV significantly outperforms existing methods, demonstrating consistent advantages across all models and benchmarks. Specifically, for LLaVA-7B, LLaVA-13B, LLaVA-OneVision, InternVL2, and Qwen2.5-VL, the average improvements relative to the base model are , , , and , respectively. Our method performs particularly well for LLaVA-OneVision on VizWiz, with an improvement of . Besides, our method appears to be more effective and robust on the VQA task, where other methods often yield negative gains. (2) Our approach shows the most significant performance advantage on MMMU compared to the API. With an average score of for our method versus for theirs, the relative difference of underscores the importance of accurate case-level visual cues retrieval instead of benchmark-level visual prompts selection. The other two visual prompting methods, lacking the ability to adapt to different queries, do not perform well on reasoning tasks. (3) Our method further supplements inherently less accurate models. For instance, LLaVA-7B equipped with our approach achieves higher accuracy than the 13B-scale model on 5 out of 8 tasks. This means that AutoV can be utilized for model efficiency on certain specific tasks: a few additional MLP layers are a better choice compared with hundreds of times larger parameter sizes. (4) Both Qwen2.5-VL and InternVL2 integrate the retrieval optimal prompting strategy from AutoV, which is pre-trained on LLaVA-OneVision. This integration results in an average accuracy gain of . Notably, the improvements are consistent and observed in all individual tasks, including complex reasoning (e.g., MathVista) and vision-centric scenarios (e.g., MMStar). This demonstrates that the effectiveness of the retrieval-enhanced prompting strategy is model-agnostic and generalizes well across heterogeneous LVLM architectures, regardless of pretraining scales or backbone variations.
Additionally, we extend the evaluation to a broader range of vision-language benchmarks, including CVBench, BLINK, Real-World QA, and MMVP, as well as classic tasks such as image captioning, visual grounding, and multimodal classification. We conduct experiments using both LLaVA-OneVision and Qwen2.5-VL as base models to examine the generalization capability of AutoV. The detailed quantitative results are provided in the Appendix D.
4.3 Analysis
4.3.1 Comparison with Other Retrieval Methods

### v3

# AutoV: Loss-Oriented Ranking for Visual Prompt Retrieval in LVLMs

@@ Abstract
Inspired by text prompts in large language models, visual prompts have been explored to enhance the perceptual capabilities of large vision-language models (LVLMs). However, performance tends to saturate under single visual prompt designs, making further prompt engineering increasingly ineffective. To address this limitation, we shift from prompt engineering to prompt retrieval and propose AutoV, a lightweight framework for instance-adaptive visual prompt identification. Given an input image and a textual query, AutoV automatically locates the most suitable visual prompt from a diverse candidate pool. Training such a retrieval framework requires prompt-level supervision, yet prompt quality is inherently ambiguous and difficult to assess reliably, even for humans. To enable automatic supervision, we evaluate visual prompts using a pre-trained LVLM and label them according to their prediction losses. Using the loss-oriented ranking as a robust training signal, AutoV learns to retrieve the query-aware optimal prompt for each instance without manual annotation. Experiments indicate that AutoV enhances the performance of various LVLMs on image understanding, captioning, grounding, and classification tasks. For example, AutoV improves LLaVA-OV by on VizWiz and boosts Qwen2.5-VL by on MMMU, respectively.
@@ 1 Introduction
Recent advancements in large language models (LLMs) [brown2020language, achiam2023gpt, bai2023qwen, bai2025qwen2, touvron2023llama, peng2023instruction, bi2024deepseek] have spurred significant progress in large vision-language models (LVLMs), which connect visual encoders with language model decoders to effectively adapt textual reasoning capabilities to visual understanding tasks [liu2023improvedllava, Qwen-VL, chen2023internvl, hurst2024gpt, li2024llava, team2024gemini, team2023gemini]. Consequently, visual input plays a crucial role in LVLMs, enabling precise perception and fine-grained grounding across various multimodal scenarios.
Inspired by textual prompting in LLMs [kojima2022large, wei2022chain, zhou2022cocoop], visual prompting has emerged as an effective paradigm for guiding LVLM attention [zhang2023vpgtrans, sun2023alphaclip, lin2024drawandunderstand, wu2024visual]. By injecting structured visual cues, such as blur masks, bounding circles, or attention heatmaps, visual prompts encourage models to focus on task-relevant regions. However, existing approaches predominantly rely on fixed, heuristically designed prompts. While these handcrafted strategies can yield improvements on certain benchmarks, they implicitly assume that a single prompt design generalizes across diverse visual scenes and textual queries.
In practice, this assumption rarely holds. As shown in Figure 1(a), heuristic prompts tend to approach benchmark-specific performance ceilings, leaving limited room for improvement through prompt engineering alone. Moreover, prompt effectiveness varies across tasks and instances. For example, in Figure 1(b), attention-based masks may benefit OCR-sensitive tasks, while highlight circles emphasize salient objects for detection. These observations suggest that the optimal visual prompt is inherently instance-dependent, making fixed prompt designs fundamentally limited. This insight motivates a paradigm shift from prompt engineering to prompt retrieval: instead of searching for a universally optimal visual prompt, we ask a more flexible question—given a specific image-query pair, which visual prompt is most suitable?
@@
To this end, we introduce AutoV. This innovative visual-prompting retrieval framework dynamically selects the optimal visual prompt from a set of candidates, explicitly tailored to the textual query and the input image. Specifically, we first generate various visual prompt candidates and encode their representations. This diversity enables the framework to capture distinct visual representations that may be optimal for various query-image pairs. Next, we design a lightweight ranking network that efficiently integrates visual and textual information to predict the ranking preferences over prompt candidates. Finally, the highest-ranking candidate is chosen as the input for LVLMs during inference.
A central challenge in learning such a retrieval framework lies in supervision. As illustrated in Figure 1(b), the quality of visual prompts is often ambiguous and task-dependent, making reliable human annotation difficult, particularly for OCR or reasoning-heavy tasks. To address this issue, we introduce a fully automated supervision strategy: each prompt candidate is evaluated by a pre-trained LVLM and ranked according to its prediction loss. This loss-oriented ranking provides a stable training signal, enabling the ranking network to learn query-aware prompt preferences without manual annotation.
Extensive experiments demonstrate that AutoV consistently enhances a wide range of LVLMs across diverse tasks, including image understanding, captioning, grounding, and classification. For example, AutoV improves LLaVA-OneVision by on VizWiz and boosts Qwen2.5-VL by on MMMU. Beyond benchmark gains, AutoV generalizes robustly across closed-source models and integrates seamlessly without additional fine-tuning of the backbone LVLM.
In summary, our contributions are threefold:
@@
We introduce AutoV, a novel automatic visual prompting framework for large vision-language models that adaptively retrieves the optimal visual prompt from a diverse candidate pool, explicitly tailored to textual query.
@@
We perform comprehensive experiments to validate the efficacy and robustness of AutoV across tasks, with consistent and significant improvements over existing visual prompt engineering. Once trained, AutoV integrates seamlessly into various LVLMs without additional fine-tuning.
@@ 2.1 Large Vision-Language Models
Recent advances in large language models [radford2019language, brown2020language, achiam2023gpt, touvron2023llama, peng2023instruction, bi2024deepseek] have catalyzed a new wave of large vision-language models [liu2023improvedllava, alayrac2022flamingo, Qwen-VL, chen2023internvl, li2024mini, li2023blip]. By coupling a high-capacity visual encoder with an LLM decoder, LVLMs transfer textual reasoning to visual understanding, and the quality of input images determines the lower bound of multimodal abilities. For instance, the popular LLaVA-OneVision [li2024llava] and Qwen2.5-VL [bai2025qwen2] adopt a higher and dynamic resolution recipe for training to enhance their fine-grained reasoning, while vision-centric scaling in InternVL2 [cai2024internlm2] enlarges the ViT [dosovitskiy2020image] backbone to boost grounding capability. However, the above methods require training times of several hundred GPU days. Here, can we further boost the existing LVLMs by optimizing the visual input itself during the inference stage? The answer is Yes. Visual prompting [sun2023alphaclip, lin2024drawandunderstand, jiang2024joint, zhang2024exploring, api, yang2023fine, shtedritski2023does] is an emerging and effective technique in this direction.
2.2 Visual Prompting for LVLMs
@@
LVLMs relying solely on textual inputs exhibited two key limitations: visual hallucinations [bai2024hallucination, huang2024visual] and language-driven biases [wu2024commit, qu2024unified]. To better mitigate these challenges, visual prompting [zhang2023vpgtrans, sun2023alphaclip, lin2024drawandunderstand, wu2024visual] has emerged as an alternative way of supplying cues in the visual modality. Compared with textual prompts, it is more concise and explicit, while allowing guidance that reaches fine-grained, pixel-level details. Early work [shtedritski2023does] shows that simply drawing a red circle around a target object steers CLIP [CLIP] attention, yielding great performance on referring-expression and keypoint-localization benchmarks. FGVP [yang2023fine] extends this idea with fine-grained prompting, replacing coarse shapes with blur-reversal prompts based on segmentation masks, further improving accuracy on RefCOCO [zhang2019referring]. AlphaCLIP [sun2023alphaclip] trains an augmented version of CLIP with an auxiliary alpha channel for indicating regions at the expense of increased training overhead. API [api] proposes attention prompting on images, overlaying text-guided attention heatmaps. Although previous studies have explored various visual prompts, the marginal gains brought by such engineering designs are gradually diminishing. To address this limitation, we propose a visual prompt retrieval mechanism that integrates existing prompt designs, exploiting their complementary advantages.
2.3 Visual Retrieval for LVLMs
@@
Existing visual prompt retrieval pipelines generally adhere to three main paradigms. (1) The heuristic paradigm identifies benchmark-level optimal parameters through extensive evaluation without instance-specific adaptation. (2) The random paradigm offers a baseline characterized by instance-level variation. (3) The Mixture-of-Experts (MoE) paradigm [fedus2022switch, bao2022vlmo, zong2024mova] provides the most direct form of adaptation by utilizing GateNet specifically for visual retrieval. Notably, AutoV differs fundamentally from retrieval-augmented methods [qi2024rora, xing2025re]. Instead of retrieving external visual data with additional finetuning, AutoV performs lightweight in-model prompt selection based solely on the input image features within a compact candidate pool, resulting in significantly improved efficiency.
@@ 3.3 Reward Loss Supervision
Automated Data Generation. Training a retrieval model requires prompt-level supervision, yet the quality of a visual prompt is inherently ambiguous and lacks a reliable evaluation criterion, even for humans. To address this, we construct supervision signals directly from the intrinsic behavior of a pre-trained LVLM. Our approach is grounded in a simple intuition: a superior visual prompt should induce lower conditional language modeling loss given the visual-question pair. Rather than relying on one-hot correctness signals, which may reflect the LVLM’s language priors instead of true visual grounding [xiao2024can, yu2024frame], we optimize the ranking network using relative losses. In this way, the loss serves as a calibrated proxy for model confidence, where lower loss implies stronger alignment between the prompt and the input instance [guo2017calibration, zhang2024unveiling].
@@ 3.4 Robust Inference Pipeline
During the inference, we input multiple candidate visual prompts along with the query into the large vision-language model and pass them through our ranking network, as shown in Figure 2. The network outputs a reward scalar for each visual prompt, and we select the highest-scoring one. Its visual tokens are concatenated with the text embeddings and decoded by the LLM to generate the final answer. Besides, to reduce distributional bias, we introduce a pre-filtering step: the prompt whose visual features are farthest from other candidates, as measured by cosine distance, is removed in advance, further mitigating the negative effects of suboptimal prompts and enhancing robustness.
@@ 4 Experiments
4.1 Experimental Setup
Model architectures. We apply our AutoV to four mainstream open-source large vision-language models, including the LLaVA-1.5 series [liu2024visual], LLaVA-OneVision [li2024llava], Qwen2.5-VL [bai2025qwen2], and InternVL2 [chen2024expanding]. In particular, LLaVA-1.5 series build on CLIP-ViT [CLIP] and Llama2 [touvron2023llama2]. LLaVA-OneVision enhances it with SigLIP [siglip] and Qwen2 [bai2023qwen]. Qwen2.5-VL introduces a dynamic-resolution encoder, and InternVL2 integrates InternViT [chen2023internvl] with InternLM2 [cai2024internlm2].
Evaluation benchmarks. AutoV is extensively evaluated on fourteen vision-language benchmarks, covering video reasoning (CVBench [tong2024cambrian]), vision-centric tests (BLINK [fu2024blink], MMVP [tong2024eyes], MMStar [chen2024we]), comprehensive multimodal evaluation (MM-Vet [yu2023mm]), knowledge/math reasoning (MMMU [yue2024mmmu], MathVista [lu2023mathvista]), robustness and VQA (VizWiz [bigham2010vizwiz], Real-World QA [RealworldQA], TextVQA [singh2019towards], LLaVA-Wild [liu2023improvedllava]), grounding (RefCOCO+ [yu2016modeling]), captioning (COCO-Caption2017 [lin2014microsoft]), and multimodal classification (ImageNet-1K [deng2009imagenet]).
| Method | MMMU | VizWiz | MMVet | VQA | LLaVA | MMStar | MathVista |
| Base | 36.3 | 50.0 | 30.6 | 58.2 | 65.4 | 33.2 | 34.2 |
| FGVP | 36.9 | 47.5 | 29.0 | 46.3 | 55.6 | 33.0 | 34.6 |
| Circle | 37.0 | 49.6 | 31.3 | 55.5 | 62.2 | 33.4 | 35.0 |
| API | 37.4 | 51.3 | 31.8 | 58.4 | 67.2 | 34.0 | 35.7 |
| AutoV | 38.7 | 52.1 | 32.9 | 60.6 | 68.6 | 35.2 | 36.9 |
| Base | 36.7 | 53.6 | 36.1 | 61.2 | 66.6 | 32.8 | 35.0 |
| FGVP | 36.9 | 46.8 | 29.7 | 49.2 | 63.7 | 31.9 | 35.3 |
| Circle | 36.9 | 53.8 | 31.2 | 58.1 | 66.9 | 33.5 | 36.2 |
| API | 37.3 | 54.6 | 37.6 | 62.1 | 67.6 | 33.0 | 37.0 |
| AutoV | 38.9 | 56.3 | 38.0 | 62.7 | 69.5 | 34.2 | 38.3 |
| Base | 49.4 | 58.2 | 48.8 | 76.1 | 80.6 | 61.8 | 63.2 |
| FGVP | 49.6 | 58.9 | 49.1 | 65.4 | 70.8 | 58.8 | 63.4 |
| Circle | 50.4 | 60.2 | 47.2 | 77.2 | 80.8 | 62.7 | 65.4 |
| API | 50.8 | 66.9 | 51.5 | 77.7 | 81.9 | 63.0 | 64.9 |
| AutoV | 54.0 | 68.4 | 52.9 | 78.0 | 85.3 | 64.5 | 66.8 |
| Base | 48.6 | 58.6 | 46.5 | 77.0 | 74.6 | 62.0 | 58.3 |
| FGVP | 49.6 | 58.3 | 44.3 | 70.7 | 75.7 | 60.0 | 58.9 |
| Circle | 48.3 | 57.5 | 45.8 | 74.9 | 75.4 | 63.4 | 58.4 |
| API | 50.5 | 61.2 | 48.6 | 77.0 | 76.2 | 63.9 | 59.4 |
| AutoV | 51.7 | 64.8 | 50.2 | 78.0 | 79.5 | 65.8 | 70.5 |
| Base | 50.1 | 69.5 | 56.6 | 83.0 | 98.0 | 62.4 | 68.2 |
| FGVP | 49.6 | 70.8 | 43.1 | 78.0 | 98.2 | 63.4 | 68.0 |
| Circle | 51.6 | 69.2 | 44.5 | 80.5 | 97.5 | 65.1 | 69.1 |
| API | 51.0 | 72.1 | 58.0 | 85.3 | 100.6 | 63.0 | 67.6 |
| AutoV | 53.9 | 74.4 | 59.1 | 86.1 | 102.3 | 65.6 | 70.2 |
@@ 4.2 Main Results
Our results are listed in Table 1, and we compare AutoV with three previous visual prompting methods designed for LVLMs, including FGVP [yang2023fine], RedCircle [shtedritski2023does], and API [api]. The training data and implementation details can be found in the Appendix A and C, respectively. Here, we employ the above three types of visual prompts as the candidate pool111Includes four prompts extracted from different transformer layers of CLIP in API [api], with two additional prompts of distinct types, resulting in six images in total. for AutoV to retrieve from optimally.
@@
The four main observations from the experimental results are as follows: (1) For LLaVA-7B, LLaVA-13B, LLaVA-OneVision, InternVL2, and Qwen2.5-VL, the average improvements of AutoV relative to the base model are , , , , and , respectively. Our method performs particularly well for LLaVA-OneVision on VizWiz, with an improvement of . (2) AutoV significantly outperforms existing visual prompt engineering methods. For instance, on MMMU, our approach achieves an average gain of , compared to for API, yielding a absolute improvement. This margin highlights the importance of accurate case-level visual cue retrieval, as opposed to benchmark-level prompt design. In contrast, other visual prompting strategies that lack query-adaptive mechanisms tend to underperform. (3) Our method further supplements inherently less accurate models. For instance, LLaVA-7B equipped with our approach achieves higher accuracy than the 13B-scale model on 4 out of 7 tasks. This means that AutoV can be utilized for model efficiency on certain specific tasks: a few additional MLP layers are a better choice compared with hundreds of times larger parameter sizes. (4) Both Qwen2.5-VL and InternVL2 directly adopt the retrieval optimal prompting strategy from AutoV, which is pre-trained on LLaVA-OneVision. This integration results in an average accuracy gain of . This demonstrates that the retrieval-enhanced prompting strategy is model-agnostic and generalizes well across heterogeneous LVLM architectures, regardless of pretraining scales or vision encoders. Additionally, we extend the evaluation to a broader range of vision-language benchmarks in Table 2. AutoV consistently enhances multimodal understanding and reasoning across diverse domains, highlighting its strong cross-task generalization in retrieval.
| Method | CVBench | BLINK | RealWorld | MMVP | COCOCap | RefCOCO+ | ImageNet-1K |
| Base | 79.6 | 48.2 | 66.0 | 60.7 | 61.1 | 68.8 | 92.3 |
| API | 80.7 | 49.4 | 66.4 | 61.5 | 61.9 | 69.4 | 92.6 |
| AutoV | 83.0 | 51.6 | 68.9 | 64.2 | 63.7 | 71.8 | 95.0 |
| Base | 80.6 | 56.4 | 68.5 | 61.3 | 40.7 | 72.6 | 94.7 |
| API | 81.1 | 57.3 | 69.2 | 61.7 | 41.4 | 73.5 | 95.4 |
| AutoV | 83.0 | 59.0 | 71.1 | 64.3 | 43.3 | 75.2 | 97.1 |
5 Analysis
5.1 Comparison with Other Retrieval Methods